1.) Two spheres of mass 1.0kg and 4.0kg are joined by a light rod so that their centers are 0.20m apart. Calculate the position of center of gravity.
1
Expert's answer
2021-10-21T16:33:30-0400

Given:

"m_1=1.0\:\rm kg"

"m_2=4.0\:\rm kg"

"l=0.20\:\rm m"


The center of gravity of the system

"x_c=\frac{x_1m_1+x_2m_2}{m_1+m_2}=\frac{0*m_1+l*m_2}{m_1+m_2}"

"x_c=\frac{0.20*4.0}{1.0+4.0}=0.16\:\rm m\quad(from\;\it m_{\rm1})"